### v1

# A New Family of Binary Sequences via Elliptic Function Fields over Finite Fields of Odd Characteristics

Abstract
Motivated by the constructions of binary sequences by utilizing the cyclic elliptic function fields over the finite field $\mathbb{F}_{2^{n}}$ by Jin et al. in [IEEE Trans. Inf. Theory 71(8), 2025], we extend the construction to the cyclic elliptic function fields with odd characteristic by using the quadratic residue map $\eta$ instead of the trace map used therein. For any cyclic elliptic function field with $q+1+t$ rational points and any positive integer $d$ with $\gcd(d,q+1+t)=1$, we construct a new family of binary sequences of length $q+1+t$, size $q^{d-1}-1$, balance upper bounded by $(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d,$ the correlation upper bounded by $(2d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+2d$ and the linear complexity lower bounded by $\frac{q+1+2t-d-(d+1)\cdot\lfloor 2\sqrt{q}\rfloor}{d+d\cdot\lfloor 2\sqrt{q}\rfloor}$ where $\lfloor x\rfloor$ stands for the integer part of $x\in\mathbb{R}$.
I Introduction
Binary sequences with low correlations (autocorrelation and cross-correlation) and higher linear complexities are widely applied in the digital communication, cryptography and signal processing.
There are several renowned constructions of binary sequences with good parameters via different methods. The oldest family of binary sequences of length $2^{n}-1$, size $2^{n}+1$ having three-level out-of-phase auto- and cross-correlation for odd integer $n$ are called Gold sequences. In 2006, a family of binary sequences of length $p$ which is an odd prime, family size (p1)/2(p-1)/2, and correlation bounded by $5+2\sqrt{p}$ was given in [16] which is called Weil sequences. The idea of constructing Weil sequences is to derive sequences from quadratic residue based on Legendre sequences using a method shift-and-add, see [1, 4]. In 2011, the modified Gold sequences were generalized and a family of binary sequences with length $2^{n}-1$ for $n=2m+1$, size $2^{\ell n}+\cdots+2^{n}-1$, and correlation $2^{m+\ell}+1$ for each $1\leq\ell\leq m$ are obtained by Zhou and Tang, see [22]. There are also some works contributing to the construction of binary sequences with length of other forms such as $2(2^{n}-1)$ and $(2^{n}-1)^{2}$, see [18]. The cyclic structure of multiplicative group $\mathbb{F}_{2^{n}}^{*}$ is also applied in the designing of the binary sequences, see [3, 19] etc. In Table I, we list some parameters of the sequence families.
TABLE I: Parameters of sequence families
Sequence Length NN Family Size Bound of Correlation
Gold(odd)([1]) $2^{n}-1$, $n$ odd $N+2$ $1+2\sqrt{2}\sqrt{N+1}$
Gold(even)([1]) $2^{n}-1$, $n=4k+2$ $N+2$ $1+2\sqrt{N+1}$
Kasami(small)([10]) $2^{n}-1$, $n$ even $\sqrt{N+1}$ $1+\sqrt{N+1}$
Kasami(large)([10]) $2^{n}-1$, $n=4k+2$ $(N+2)\sqrt{N+1}$ $1+2\sqrt{N+1}$
Gong([3]) $(2^{n}-1)^{2}$, $2^{n}-1$ prime $\sqrt{N}$ $3+2\sqrt{N}$
Tang et al. ([18]) $2(2^{n}-1)$, $n$ odd $\frac{N}{2}+1$ $2+\sqrt{N+2}$
Jin et al. ([9]) $2^{n}$ $N^{2}-1$ $1+14\sqrt{N}$
Jin et al.([9]) $\pm\sqrt{q},q=4^{n} q1q-1$ $11\sqrt{q}$
Jin et al.([9]) $\pm\sqrt{2q},q=2\times 4^{n} q1q-1$ $(10+\sqrt{2})\sqrt{q}$
$\mathbf{Our}$ $\mathbf{construction}$ $q+1\pm\sqrt{3q},q=3^{n}$, $n$ odd q1q-1 $4+(10+\sqrt{3})\sqrt{q}$
$\mathbf{Our}$ $\mathbf{construction}$ $q+1\pm\sqrt{q},q=5^{n}$, $n$ even q1q-1 $4+11\sqrt{q}$
The theory of function fields has been widely applied in the designing of the sequences. Now we recall some constructions based on the structures of some function fields. In [7, 8], the cyclic structure of all rational points of rational function fields are employed to construct a family of binary sequences of length $2^{n}+1$ ($p^{n}+1$), size $2^{n}-1$ ($p^{n}-1$) and correlation bounded by $2^{(n+2)/2}$ ($4+\lfloor 2\cdot p^{n/2}\rfloor$) via cyclotomic function fields. The low-correlation, large linear span sequences are constructed by utilizing the theory of Artin-Schreier extensions of function fields and in particular, Xing et al. gave the general construction on the function fields over the finite fields with even characteristic, see [20]. In [5], Hu et al. presented a new construction by utilizing the theory of Kummer extensions of function fields and they extended the results of Xing et al.. Luo et al. constructed sequences from rational function fields and cyclotomic function fields with high nonlinear complexity in [11]. Hu et al. studied a class of $p$-ary pseudorandom sequences from elliptic curves over the finite fields are constructed in [6], and they investigated their periods, linear complexities, aperiodic correlation etc. While several constructions of binary sequences with favorable parameters have been introduced, the available choices for sequence lengths remain limited. Recently, Jin et al. constructed a novel family of binary sequences which is constructed on rational points on the cyclic elliptic curves over finite fields with even characteristics in [9]. In this paper, we construct a family of sequences over the elliptic curves over the finite fields with odd characteristic by using the theory of Kummer extensions of elliptic function fields and the quadratic residue map, which extends the construction by Jin et al. in [9] to the cyclic elliptic function fields over the finite fields with odd characteristic. Then we obtain a new family of binary sequences and give the upper-bound for the correlation and the lower bound for the linear complexity.
I-A Our Main Results and Techniques
Suppose $E/\mathbb{F}_{q}$ is an elliptic function field and $E(\mathbb{F}_{q})$ is denoted by the set of the $\mathbb{F}_{q}$-rational places of $E/\mathbb{F}_{q}$. It is well-known that $E(\mathbb{F}_{q})$ has an abelian group structure where we denote by $\oplus$ the group operation, see [17]. For any $P\in E(\mathbb{F}_{q})$ and $j\in\mathbb{Z}$, $[j]P$ means the addition of $j$ times of $P$ i.e.,
$j]P=\underbrace{P\oplus\cdots\oplus P}_{j}$
If $(E(\mathbb{F}_{q}),\oplus)$ form a cyclic group, then $E/\mathbb{F}_{q}$ is called a cyclic elliptic function field.
In the following, we shall focus on the cyclic elliptic function fields over finite fields with odd characteristic. Let $E/\mathbb{F}_{q}$ be a cyclic elliptic function field with $q+1+t$ rational places and $d\geq 2$ be a positive integer with $\gcd(d,q+1+t)=1$, we construct a family of binary sequences of length $q+1+t$, size $q^{d-1}-1$, linear complexity lower bounded by
$\frac{q+1+2t-d-(d+1)\cdot\lfloor 2\sqrt{q}\rfloor}{d+d\cdot\lfloor 2\sqrt{q}\rfloor}$
where $\lfloor x\rfloor$ stands for the integer part of $x\in\mathbb{R}$ and correlation upper bounded by
$(2d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+2d.$
By taking explicit cyclic elliptic function fields, we extend the families of the binary sequences with new parameters, see Table I.
Now we overview the techniques which shall be used throughout this paper. The automorphism group of an elliptic function field $E$ over $\mathbb{F}_{q}$ is in the form of the semi-product $T_{E}\rtimes\mathrm{Aut}(E,O)$, where $\mathrm{Aut}(E,O)$ consists of the $\mathbb{F}_{q}$-automorphisms of $E$ fixing the infinity place $O$, and $T_{E}$ corresponds to the translation group $E$. Let $Q$ be a place of $E$ with $\deg(Q)=d\geq 2$ such that $\gcd(d,q+1+t)=1$ and by the Riemann-Roch theorem, we have $\ell(Q)=\dim(\mathcal{L}(Q))=\deg(Q)+1-g(E)=\deg(Q)=d$ because of $\deg(Q)=d\geq 2g(E)-1=1$. Obviously, we have the decomposition $\mathcal{L}(Q)=\mathbb{F}_{q}\oplus V$. The vector space $V$ is called the complementary dual space of $\mathbb{F}_{q}$ in $\mathcal{L}(Q)$. The elements in $V\setminus\{0\}$ can be listed as $\{z_{1},z_{2},\cdots,z_{q^{d-1}-1}\}$.
Now we define the quadratic residue map $\eta:\mathbb{F}_{q}\rightarrow\mathbb{F}_{2}$ with $q$ an odd prime power by
$\eta(a)=\begin{cases}0,\ \text{if there exists}\ b\in\mathbb{F}_{q}\ \text{such that}\ a=b^{2}\ \text{or}\ a=0\\ 1,\ \text{otherwise}.\end{cases}$
It is clear that $\eta(a_{1}a_{2})=\eta(a_{1})+\eta(a_{2}),$ for any $a_{1},a_{2}\in\mathbb{F}_{q}^{*}$.
Let $\sigma_{P}$ be a generator of the cyclic translation group $T_{E}$ for some rational place $P$ on $E(\mathbb{F}_{q})$. That is, $P_{j}=\sigma^{j}_{P}[O]=[j]P$ for $0\leq j\leq q+t$ are all rational places of $E$. For any function $z_{i}\in V\setminus\{0\}$ with $1\leq i\leq q^{d-1}-1$, the binary sequence $\mathbf{s}_{i}$ associated to $z_{i}$ is defined as follows:
$\mathbf{s}_{i}=(s_{i,0},s_{i,1},\cdots,s_{i,q+t})\ \text{with}\ s_{i,j}=\eta(z_{i}(P_{j}))\ \text{for}\ 0\leq j\leq q+t.$
Then the family of binary sequences $\mathbf{s}_{i}$ has length $q+1+t$ and size $q^{d-1}-1$. The problems of computing the correlation and linear complexity are transformed into counting the number of rational places of certain Kummer extensions $E_{j,k}=E(y)$ where $y^{2}=z_{j}\cdot\tau^{-u}(z_{k})$ for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$ by the cyclic structure of $T_{E}$.
By contrast, Jin et al. in [8] constructed binary sequences from the finite field $\mathbb{F}_{2^{n}}$ by utilizing the trace map $\mathrm{Tr}:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2}$ where the problems of computing the correlation and linear complexity are transformed into counting the number of rational places of the Artin-Schreier extensions $E^{\prime}_{j,k}=E(y)$ where $y^{2}+y=z_{j}+\tau^{-u}(z_{k})$ for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$.
Although our bound applies to more general cases of the finite fields, both the zero places and pole places of $z_{j}\cdot\tau^{-u}(z_{k})$ are ramified in the Kummer extensions while only the pole places of $z_{j}+\tau^{-u}(z_{k})$ are ramified in the Artin-Schreier extensions for $1\leq j,k\leq q^{d-1}-1,0\leq u\leq q+t$, which induce the differences between the two bounds. But the two bounds are closed and the actual correlation is much smaller than the theoretical bound as is shown in the numerical experiments.
For cryptographic applications, we explored the property of balance. Also by counting the rational places over the Kummer extensions, we have found that the theoretical value of the balance is upper bounded by
$(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.$
By utilizing the software SageMath [2], it is also shown in table II that the actual difference is also much smaller than this theoretical bound.
I-B Organization of This Paper
The rest of this paper is organized as follows: In Section II, we provide preliminaries on binary sequences, extension theory of function fields, elliptic function fields and Kummer extensions of the function fields. In Section III, we present a theoretical construction of binary sequences with low correlation, favorable balance and large linear complexity via cyclic elliptic function fields over finite fields with odd characteristic by utilizing the theory of the Kummer extensions and we also present some examples to specify our construction. In Section IV, we give a conclusion and some possible future work on the constructions of binary sequences.
II Preliminaries
In this paper, we list some preliminaries on binary sequences, extension theory of function fields, elliptic function fields and Kummer extensions.
II-A Autocorrelation and Cross Correlation of the Binary Sequences
For a positive integer $N$, a binary sequence of length $N$ is a vector $\mathbf{s}\in\mathbb{F}^{N}_{2}$. Let $\mathcal{S}$ be a family of binary sequences in $\mathbb{F}_{2}^{N}$. For a binary sequence $\mathbf{s}=(s_{0},s_{1},\cdots,s_{N-1})$, its autocorrelation $A_{t}(\mathbf{s})$ at delay $t$ with $0\leq t\leq N-1$ is defined as
$A_{t}(\mathbf{s}):=\sum^{N-1}_{j=0}(-1)^{s_{j}+s_{j+t}},$
where $j+t$ is identified with the least non-negative integer modulo $N$. For two distinct binary sequences $\mathbf{u}=(u_{0},u_{1},\cdots,u_{N-1})$ and $\mathbf{v}=(v_{0},v_{1},\cdots,v_{N-1})$, their cross correlation at delay $t$ with $0\leq t\leq N-1$ is defined as
$C_{t}(\mathbf{u,\mathbf{v}}):=\sum^{N-1}_{j=0}(-1)^{u_{j}+v_{j+t}}.$
The correlation of the family of binary sequences $\mathcal{S}$ is defined as $Cor(\mathcal{S}):=\max\{\{|A_{t}(\mathbf{s})|:\mathbf{s}\in\mathcal{S},1\leq t\leq N-1\}\cup\{|C_{t}(\mathbf{u,v})|:\mathbf{u}\neq\mathbf{v}\in\mathcal{S},0\leq t\leq N-1\}\}$.
II-B Linear Complexity of Binary Sequences
Now we introduce the definition of the linear complexity of binary sequences. For any sequence $\mathbf{0}\neq\mathbf{s}\in\mathcal{S}$, the linear complexity of $\mathbf{s}$, denoted by $\ell(\mathbf{s})$, is the smallest positive integer $\ell$ such that there exist $\ell+1$ elements $\lambda_{0},\lambda_{1},\cdots,\lambda_{\ell}\in\mathbb{F}_{2}$ with $\lambda_{0}=\lambda_{\ell+1}=1$ satisfying
$\sum^{\ell}_{i=0}\lambda_{i}s_{i+u}=0$
for all $0\leq u\leq N-1$. Define the linear complexity of the family of sequences $\mathcal{S}$ by
$\mathrm{LC}(\mathcal{S}):=\min\{\ell(\mathbf{s}):\mathbf{s}\in\mathcal{S}\}.$
II-C Extension Theory of Function Fields
Let $F/\mathbb{F}_{q}$ be an algebraic function field with genus $g(F)$ over the constant field $\mathbb{F}_{q}$. Any place of $F$ which is invariant under the action of Galois group $Gal(\overline{F}_{q}/\mathbb{F}_{q})$ is called rational. Denote by $N(F)$ the number of all the rational places of $F$. The Serre bound improves the well-known Hasse-Weil bound and it will be used to the estimation of the correlation and linear complexity in the followings:
$|N(F)-q-1|\leq g(F)\cdot\lfloor 2\sqrt{q}\rfloor.$
Here $\lfloor x\rfloor$ stands for the integer part of $x\in\mathbb{R}$.
A divisor is a formal sum of places. Denote by $\mathbb{P}_{F}$ and $\mathbb{D}_{F}$ the set of all places and divisors respectively. Let $v_{P}$ be a normalized discrete valuation of $F$. For divisor $G=\sum_{P\in\mathbb{P}_{F}}v_{P}(G)P\in\mathbb{D}_{F}$, the support of $G$ is denoted by $\mathrm{Supp}(G)=\{P\in\mathbb{P}_{F}\ |\ v_{P}(G)\neq 0\}$. The degree of divisor $\deg(G)$ is defined as $\deg(G)=\sum_{P\in\mathbb{P}_{F}}v_{P}(G)\deg(P)$. A divisor $G=\sum_{P\in\mathbb{P}_{F}}v_{P}(G)P$ is called effective if $v_{P}(G)\geq 0$ for all $P\in\mathrm{Supp}(G)$.
For any rational function $f\in F\setminus\{0\}$, denote by $(f)$, $(f)_{0}$ and $(f)_{\infty}$ principal divisor, zero divisor and pole divisor respectively. The Riemann-Roch space associated to a non-negative divisor $G$ is given by the following:
$\mathcal{L}(G):=\{f\in F\setminus\{0\}\ |\ (f)+G\geq 0\}\cup\{0\}$
and it is a finite $\mathbb{F}_{q}$-dimensional vector space, and we denote by $\ell(G)$ the dimension of $\mathcal{L}(G)$. If $\deg(G)\geq 2g(F)-1$, then $\ell(G)=\deg(G)+1-g(F)$ from the well-known Riemann-Roch theorem and details can be checked at [13].
Denote by $\mathrm{Aut}(F/\mathbb{F}_{q})$ the automorphism group of $F$ over $\mathbb{F}_{q}$, i.e.,
$\mathrm{Aut}(F/\mathbb{F}_{q})=\{\sigma:F\rightarrow F\mid\sigma\ \text{is an $\mathbb{F}_{q}-$automorphism of}\ F\}.$
In the following, we shall consider the group action of automorphism group $\mathrm{Aut}(F/\mathbb{F}_{q})$ on the set of places of $F$. For any place $P$ of $F$, $\sigma(P)$ is a place of $F$ for any automorphism $\sigma\in\mathrm{Aut}(F/\mathbb{F}_{q})$ and we have the following results.
Lemma II.1.
(see [14]) For any automorphism $\sigma\in\mathrm{Aut}(F/\mathbb{F}_{q})$, $P\in\mathbb{P}_{F}$ and $f\in F$, we have
  1. 1.
    $\deg(\sigma(P))=\deg(P)$;
  2. 2.
    $v_{\sigma(P)}(\sigma(f))=v_{P}(f)$;
  3. 3.
    $\sigma(f)(\sigma(P))=f(P)$, provided that $v_{P}(f)\geq 0$.
Let $T/\mathbb{F}_{q}$ be the finite extension of $F/\mathbb{F}_{q}$. Denote by $\mathrm{Diff}(T/F)$ the different of $T/F$. For any $P\in\mathbb{P}_{F}$ and $Q\in\mathbb{P}_{T}$ with $Q|P$, let $d(Q|P),e(Q|P),f(Q|P)$ be the different exponent, ramification index and relative degree of $Q|P$, respectively. The different of $T/F$ is defined as $\mathrm{Diff}(T/F)=\sum_{P\in\mathbb{P}_{F}}\sum_{Q|P}d(Q|P)Q$ and different exponents can be determined by Dedekind’s Different Theorem or Hilbert’s Different Theorem by considering the orders of the higher ramification groups and whether $e(Q|P)$ is relatively prime to $p$, see [13]. Then the Hurwitz genus formula (see [13]) yields
$2g(T)-2=[T:F]\cdot(2g(F)-2)+\deg(\mathrm{Diff}(T/F)).$
II-D Elliptic Function Fields
An elliptic function field $E/\mathbb{F}_{q}$ is an algebraic function field over $\mathbb{F}_{q}$ of genus one. Two divisors $A,B\in\mathbb{D}_{E}$ are called equivalent if there exists $z\in E^{*}$ such that $A=B+(z)$, which is denoted by $A\sim B$. The details can be checked in [17].
The two elliptic curves $E_{1}$ and $E_{2}$ are isogenous if there is a non-constant smooth $\mathbb{F}_{q}$-morphism that sends the zero of $E_{1}$ to the zero of $E_{2}$ and it is also well known that two elliptic curves over $\mathbb{F}_{q}$ are isogenous if and only if they have the same number of rational points and details can be checked in [12].
Lemma II.2.
The isogeny classes of elliptic curves over $\mathbb{F}_{q}$ for $q=p^{n}$ are in one-to-one correspondence with the rational integers $t$ satisfying one of the following conditions:
  1. 1.
    $|t|\leq 2\sqrt{q}$ and $\gcd(t,p)=1$;
  2. 2.
    $t=\pm 2\sqrt{q}$ if $n$ is even;
  3. 3.
    $t=\pm\sqrt{q}$ if $n$ is even and $p\not\equiv 1\mod 3$;
  4. 4.
    $t=\pm p^{(n+1)/2}$ if $n$ is odd and $p=2$ or $3$;
  5. 5.
    $t=0$ if either $1)$ $n$ is odd or $2)$ $n$ is even and $p\not\equiv 1\mod 4$.
Furthermore, an elliptic curve in the isogeny class corresponding to $t$ has $q+1+t$ rational places. The following lemma reveals the structure of the additive group of rational points on the elliptic curve and details can be checked in [15].
Lemma II.3.
Let $\mathbb{F}_{q}$ be the finite field with $q=p^{n}$ elements. Let $h=\prod_{\ell}\ell^{h_{\ell}}$ be a possible number of rational places of an elliptic function field $E$ of $\mathbb{F}_{q}$. Then all the possible groups $E(\mathbb{F}_{q})$ are of the following form
$\mathbb{Z}/p^{h_{p}}\mathbb{Z}\times\prod_{\ell\neq p}(\mathbb{Z}/p^{a_{\ell}}\mathbb{Z}\times\mathbb{Z}/p^{h_{\ell}-a_{\ell}}\mathbb{Z})$
with
  1. 1.
    In case $2)$ of Lemma II.2: each $a_{\ell}$ is equal to $h_{\ell}/2$, i.e., $E(\mathbb{F}_{q})\simeq\mathbb{Z}/(\sqrt{q}\pm 1)\mathbb{Z}\times\mathbb{Z}/(\sqrt{q}\pm 1)\mathbb{Z}$.
  2. 2.
    In other cases of Lemma II.2: $a_{\ell}$ is an arbitrary integer satisfying $0\leq a_{\ell}\leq\min\{v_{\ell}(q-1),\lfloor h_{\ell}/2\rfloor\}$. In cases $3)$ and $4)$ of Lemma II.2: $E(\mathbb{F}_{q})\simeq\mathbb{Z}/h\mathbb{Z}$. In cases $5)$ of Lemma II.2: if $q\not\equiv-1\mod 4$, then $E(\mathbb{F}_{q})\simeq\mathbb{Z}/(q+1)\mathbb{Z};$ otherwise, $E(\mathbb{F}_{q})\simeq\mathbb{Z}/(q+1)\mathbb{Z}$ or $E(\mathbb{F}_{q})\simeq\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/\frac{q+1}{2}\mathbb{Z}.$
Let $\sigma_{P}$ be the automorphism of $E$ induced by $\sigma_{P}(Q)=P\oplus Q$ for any rational places $P$ and $Q$ of $E$, $T_{E}=\{\sigma_{P}:P\in E(\mathbb{F}_{q})\}$ be the translation group of $E$ and $\mathrm{Aut}(E,O)$ be the automorphism group of $E$ fixing the infinity place $O$ as before. The automorphism group of $E/\mathbb{F}_{q}$ is in the form $T_{E}\rtimes\mathrm{Aut}(E,O)$ and details can be checked in [12]. Then it suffices to consider the actions of the translation group $T_{E}$ at the places of $E$.
Let $F$ be the fixed subfield of $E$ with respect to the translation group $T_{E}$, i.e.,
$F=E^{T_{E}}=\{z\in E:\tau(z)=z\ \text{for any}\ \tau\in T_{E}\},$
where $\tau(z)(P)=z(\tau^{-1}(P))$ for any $P\in E(\mathbb{F}_{q})$ and $\tau\in T_{E}$. From the Galois theory, $E/F$ is a Galois extension with the Galois group $\mathrm{Gal}(E/F)\simeq T_{E}$.
Based on Lemma II.2 and Lemma II.3, we have the following corollary:
Corollary II.4.
Let $q=p^{n}$. Let $t$ be an integer satisfying one of the following conditions:
  1. 1.
    $|t|\leq 2\sqrt{q}$ and $\gcd(t,p)=1$;
  2. 2.
    $t=0$, $n$ is odd or $n$ is even with $q\not\equiv-1\mod 4$;
  3. 3.
    $t=\pm\sqrt{q}$ if $n$ is even and $p\not\equiv 1\mod 3$; $t=\pm p^{(n+1)/2}$ if $n$ is odd and $p=3$.
Then there exists a cyclic elliptic function field $E/\mathbb{F}_{q}$ with $q+1+t$ rational places.
In the following, we assume that $t$ satisfies one of the three conditions given above. Hence, there is a cyclic elliptic function field $E/\mathbb{F}_{q}$ with $q+1+t$ rational places. Then the translation group $T_{E}$ of $E$ is a cyclic group with the generator $\sigma_{P}$ and the order $q+1+t$. Let $P_{j}=\sigma_{P}^{j}(O)=[j]P$ be the rational place of $E$ for each $0\leq j\leq q+t$ and the set $\{P,P_{1},\cdots,P_{q+t}\}$ consists of all the rational places of $E$.
Let $d$ be a positive integer which is relatively prime with $q+1+t$. Then the following Lemmas illustrate that there are exactly $r=B_{d}/(q+1+t)$ places of $F=E^{T_{E}}$ of degree $d$ which split completely in the extension $E/F$.
Lemma II.5 (see [9]).
[…]
$B_{d}=\frac{1}{d}\cdot\sum_{r\mid d}\mu\left(\frac{d}{r}\right)\cdot\left(q^{r}+1-S_{r}\right),$
where
$S_{r}=\sum^{\lfloor r/2\rfloor}_{i=0}(-1)^{r-i}\frac{(r-i-1)!\cdot r}{(r-2i)!\cdot i!}t^{r-2i}q^{i}.$
In particular, we have $B_{2}=(q^{2}+q-t^{2}-t)/2$ and $B_{3}=(q^{3}-q+t^{3}-3qt-t)/3$.
Lemma II.6 (see [20, 5]).
Let $E$ be an cyclic elliptic function field. Let $P$ be a generator of $E(\mathbb{F}_{q})$. Then for any place $Q$ in $E(\mathbb{F}_{q})$ of degree $d$ with $\gcd(d,q+1+t)=1$, and any fixed positive integer $i$, the places $\sigma_{P}^{i}(Q),\sigma_{P}^{i+1}(Q),\cdots,\sigma^{i+q+t}_{P}(Q)$ are pairwise different.
In this paper, we shall utilize the quadratic residue map $\eta$ and this structure of the rational points over the cyclic elliptic function field with odd characteristic to construct a new family of binary sequences with good parameters.
II-E Kummer Extensions of Function Fields
In this subsection, we recall some properties of Kummer extensions of function fields and details can be checked at [17].
Proposition II.7.
Let $F/K$ be an algebraic function field where $K$ contains a primitive $m$-th root of unity. Suppose that $u\in F$ is an element satisfying
$u\neq w^{d}\ \text{for all}\ w\in F\ \text{and}\ d\mid m,d>1.$
Let
$F^{\prime}=F(y)\ \text{with}\ y^{m}=u.$
Such an extension $F^{\prime}/F$ is said to be a Kummer extension of $F$. We have:
  1. 1.
    The polynomial $\Phi(T)=T^{m}-u$ is the minimal polynomial of $y$ over $F$ (in particular, it is irreducible over $F$). The extension $F^{\prime}/F$ is Galois of degree $F^{\prime}:F]=m$; its Galois group is cyclic, and the automorphisms of $F^{\prime}/F$ are given by $\sigma(y)=\zeta y$, where $\zeta\in K$ is an $m$-th root of unity.
  2. 2.
    Let $P\in\mathbb{P}_{F}$ and $P^{\prime}\in\mathbb{P}_{F^{\prime}}$ be an extension of $P$. Then
    $e(P^{\prime}|P)=\frac{m}{r_{P}}\ \text{and}\ d(P^{\prime}|P)=\frac{m}{r_{P}}-1,$
    where $r_{P}:=\gcd(m,v_{P}(u))>0$ is the greatest common divisor of $m$ and $v_{P}(u)$.
  3. 3.
    Assume that there is $Q\in\mathbb{P}_{F}$ such that $\gcd(v_{Q}(u),m)=1$. Then $K$ is the full constant field of $F^{\prime}$, the extension $F^{\prime}/F$ is cyclic of order $m$, and the genus of $F^{\prime}$ satisfies
    $2g(F^{\prime})-2=m(2g(F)-2)+\sum_{P\in\mathbb{P}_{F}}(m-r_{P})\deg P.$
III New Binary sequences via Cyclic elliptic function fields with odd characteristics
Motivated by the construction of the families of the binary sequences with low correlation via the cyclotomic function fields and the structure of cyclic elliptic curves over the finite fields with even characteristic, see [7, 8, 9] respectively, we extend the constructions to the cyclic elliptic curves over the finite fields with odd characteristic.
III-A A New Construction of Binary Sequences via Cyclic Elliptic Function Fields
Let $\mathbb{F}_{q}$ be the finite field with $q=p^{n}$ and odd $p$. In this subsection, we shall utilize the cyclic group structure of rational places of cyclic elliptic function fields over $\mathbb{F}_{q}$ to construct a new family of binary sequences with low correlation and large linear complexity by utilizing the quadratic residue map $\eta$.
Let $Q$ be a place of $E$ with $\deg(Q)=d\geq 2$ such that $\gcd(d,q+1+t)=1$. The existence of such a place $Q$ can be guaranteed by Lemma II.5. Then we have the decomposition $\mathcal{L}(Q)=\mathbb{F}_{q}\oplus V$ and the vector space $V$ is called the complementary dual space of $\mathbb{F}_{q}$ in $\mathcal{L}(Q)$ and all the elements of $V\setminus\{0\}$ can be listed as $\{z_{1},z_{2},\cdots,z_{q^{d-1}-1}\}$. Let $\eta$ be the quadratic residue map defined before and $\sigma_{P}$ be a generator of the cyclic translation group $T_{E}$ and $P_{j}=\sigma^{j}_{P}[O]=[j]P$ for $0\leq j\leq q+t$ be all the rational places of $E$. Then the binary sequence $\mathbf{s}_{i}$ associated to $z_{i}$ is defined as follows:
$\mathbf{s}_{i}=(s_{i,0},s_{i,1},\cdots,s_{i,q+t})\ \text{with}\ s_{i,j}=\eta(z_{i}(P_{j}))$
for $1\leq i\leq q^{d-1}-1$ and $0\leq j\leq q+t$. Denote by $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ the family of sequences we constructed above. Now we shall show that this family of binary sequences $\mathcal{S}$ has low correlation and large linear complexity.
III-B The Balance and Correlation
In this subsection, we shall show that the family of binary sequences $\mathcal{S}=\{\mathbf{s}:1\leq i\leq q^{d-1}-1\}$ constructed above has favorable balance and low correlation.
Lemma III.1.
For any element $z_{1},z_{2}\in V\setminus\{0\}$, one has
  1. 1.
    $(z_{1}\sigma(z_{2}))_{\infty}=Q+\sigma(Q)$ for any automorphism $\sigma\in\mathrm{Gal}(E/F);$
  2. 2.
    $\deg((z_{1}\sigma(z_{2}))_{\infty})=2d$ for any automorphism $\sigma\in\mathrm{Gal}(E/F).$
Proof.
For any rational functions $z_{1},z_{2}\in V\setminus\{0\}$, the pole divisors are given by $(z_{1})_{\infty}=(z_{2})_{\infty}=Q$. By Lemma II.1, we have $v_{\sigma(Q)}(\sigma(z_{2}))=v_{Q}(z_{2})=-1$ and $(\sigma(z_{2}))_{\infty}=\sigma(Q)$. Then for any $\sigma\in\mathrm{Gal}(E/F)\setminus\{\mathrm{id}\}$we have $z_{1}\sigma(z_{2})\in\mathcal{L}(Q+\sigma(Q))$ and $\sigma(Q)\neq Q$. If $\sigma=id$, then we have $(z_{1}\sigma(z_{2}))_{\infty}=(z_{1}z_{2})_{\infty}=2Q$. Therefore, $(z_{1}\sigma(z_{2}))_{\infty}=Q+\sigma(Q)$ for any automorphism $\sigma\in\mathrm{Gal}(E/F)$.
For the second claim, for any $z_{1},z_{2}\in V\setminus\{0\}$, we have $(z_{1}z_{2})_{\infty}=2Q$ since $(z_{1})_{\infty}=(z_{2})_{\infty}=Q$ and by the relation $\deg(Q)=\deg(\sigma(Q))=d$ for $id\neq\sigma\in\mathrm{Gal}(E/F)$, we have $\deg((z_{1}\sigma(z_{2}))_{\infty})=2d$ for any automorphism $\sigma\in\mathrm{Gal}(E/F)$. ∎
Lemma III.2.
Let $\mathbb{F}_{q}$ be a finite field with $q$ an odd prime power and $E/\mathbb{F}_{q}$ an elliptic function field. Consider the Kummer extension $E(y)/E$ where $y^{2}=z$ for some function $z\in V\setminus\{0\}$. Then we have the following results:
  1. 1.
    The ramified places in $\mathbb{P}_{E}$ are those $P\in\mathbb{P}_{E}$ satisfying $\gcd(2,v_{P}(z))=1$.
  2. 2.
    The genus $g(E^{\prime})$ of the Kummer extension $E^{\prime}$ is upper bounded by
    $g(E^{\prime})\leq 1+d.$ (1)
Proof.
For the first claim, by the Proposition II.7, if any place $P\in\mathbb{P}_{E}$ is ramified in $E^{\prime}/E$ and $P^{\prime}|P$ with $P^{\prime}\in\mathbb{P}_{E^{\prime}}$, they satisfies
$e(P^{\prime}|P)=\frac{2}{r_{P}}\ \text{and}\ d(P^{\prime}|P)=\frac{2}{r_{P}}-1$
where $r_{P}=\gcd(2,v_{P}(z))$ and we have the following classifications.
If $r_{P}=\gcd(2,v_{P}(z))=1$, then the ramification index and different exponent are given by
$e(P^{\prime}|P)=2\ \text{and}\ d(P^{\prime}|P)=1.$
If $r_{P}=\gcd(2,v_{P}(z))=2$, then we have
$e(P^{\prime}|P)=1\ \text{and}\ d(P^{\prime}|P)=0$
which means $P\in\mathbb{P}_{E}$ is unramified in the Kummer extension $E^{\prime}/E$.
For the second claim, based on the analysis of the ramified places, we have determined all the ramified places in $\mathbb{P}_{E}$ in the Kummer extension $E^{\prime}/E$. According to the Hurwitz genus formula, we have the following estimation
$\begin{split}2g(E^{\prime})-2&=[E^{\prime}:E](2g(E)-2)+\mathrm{Diff}(E^{\prime}/E)\\ &=\mathrm{Diff}(E^{\prime}/E)\ (\text{by the genus}\ g(E)=1)\\ &=\deg(Q)+\sum_{P\in\mathcal{Z}(z),\ r_{P}=1}\deg P\\ &\leq d+\sum_{P\in\mathcal{Z}(z)}\deg P\\ &\leq d+\sum_{P\in\mathcal{Z}(z)}v_{P}(z)\deg P=2d.\end{split}$
where $\mathcal{Z}(z)$ is denoted by the set of the zero places of $z$.
Then we have $2g(E^{\prime})-2\leq 2d$ which is equivalent to $g(E^{\prime})\leq 1+d$ and we have the desired results. ∎
Lemma III.3 (see [20]).
Let $z\in V\setminus\{0\}$. Put $P_{i}=\tau^{i}(P)$ for all $1\leq i\leq n-1$ and $\tau\in\mathrm{Aut}(E/\mathbb{F}_{q})$, then $P_{j+n}=P_{j}$ for all $j\in\mathbb{Z}$ and $P_{\ell},P_{\ell+1},\cdots,P_{\ell+n-1}$ are $n$ pairwise distinct rational places for any fixed $\ell\in\mathbb{Z}$ and we have $z(P_{\ell+u})=z(\tau^{u}(P_{\ell}))=\tau^{-u}(z)(P_{\ell})$ for any $u<n$.
Based on the lemmas above, we have the following result of the correlation and the balance for this new family of binary sequences $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ constructed above.
First, for cryptographic applications, a binary sequence must be balanced, i.e. the number of $1$s and 0s are closed. Now we give a formal definition for this property.
Definition III.4.
Denoted by $N^{\prime\prime}_{0}(\mathbf{s})$ and $N^{\prime\prime}_{1}(\mathbf{s})$ the number of 0s and $1$s in a binary sequence $\mathbf{s}$ respectively. The binary sequence $\mathbf{s}$ is called balanced if
$|N^{\prime\prime}_{1}(\mathbf{s})-N^{\prime\prime}_{0}(\mathbf{s})|=|2N^{\prime\prime}_{1}(\mathbf{s})-n|\leq 1$
where $n$ is the length of the sequence $\mathbf{s}$.
Now we define $\Delta(\mathcal{S})\triangleq\max\{|N^{\prime\prime}_{1}(\mathbf{s})-N^{\prime\prime}_{0}(\mathbf{s})|,\ \mathbf{s}\in\mathcal{S}\}$ and $\Delta(\mathcal{S})$ is called the balance. In this subsection, we study the property of balance and it turns out that the problem is transformed into the analysis of the number of rational places for the Kummer extensions.
Theorem III.5.
Let $q=p^{n}$ for an odd prime $p$ and a positive integer $n$. Let $E/\mathbb{F}_{q}$ be a cyclic elliptic function field defined over $\mathbb{F}_{q}$ with $q+1+t$ rational places. Let $F$ be the fixed subfield of $E$ with respect to its translation group $T_{E}$ generated by $\sigma_{P}$. Let $Q$ be a place of $E$ with $\deg(Q)=d\geq 2$ such that $\gcd(d,q+1+t)=1$. Then the family of binary sequences $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ constructed above is with length $q+1+t$, size $q^{d-1}-1$ and the following inequality holds
$\begin{split}\Delta(\mathcal{S})&=\left|2N_{1}^{\prime\prime}(\mathcal{S})-(q+1+t)\right|\\ &\leq(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.\end{split}$ (2)
where $\Delta(\mathcal{S})=\max\{|N^{\prime\prime}_{1}(\mathbf{s}_{i})-N^{\prime\prime}_{0}(\mathbf{s}_{i})|,1\leq i\leq q^{d-1}-1\}$.
Proof.
Consider the Kummer extension $E_{i}^{\prime}=E(y)$ with the defining equation
$E^{\prime}_{i}:y^{2}=z_{i}$
for each $1\leq i\leq q^{d-1}-1$. By Lemma 1, we have $g(E^{\prime})\leq 1+d$.
Denote by $N^{\prime\prime}_{0}(\mathbf{s}_{i})$ the cardinality of the set $\mathcal{S}_{0}^{\prime\prime}=\{0\leq j\leq q+t:\eta(z_{i}(P_{j}))=0\}$ and $N^{\prime\prime}_{1}(\mathbf{s}_{i})$ the cardinality of the set $\mathcal{S}_{1}^{\prime\prime}=\{0\leq j\leq q+t:\eta(z_{i}(P_{j}))=1\}$ for each $1\leq i\leq q^{d-1}-1$. By the relation $N^{\prime\prime}_{0}(\mathbf{s}_{i})+N^{\prime\prime}_{1}(\mathbf{s}_{i})=q+1+t$, we have
$\begin{split}|N^{\prime\prime}_{0}(\mathbf{s}_{i})-N^{\prime\prime}_{1}(\mathbf{s}_{i})|&=|2N^{\prime\prime}_{1}(\mathbf{s}_{i})-(q+1+t)|\\ &=|2N^{\prime\prime}_{0}(\mathbf{s}_{i})-(q+1+t)|\end{split}$
then we only need to consider the cardinality $N^{\prime\prime}_{0}(\mathbf{s}_{i})$.
If $j\in\mathcal{S}^{\prime\prime}_{0}$, then $P_{j}$ splits into 2 rational places in $E^{\prime}_{i}$ or there is only one rational place lying over $P_{j}$ in $E^{\prime}_{i}$ by the definition of quadratic residue map $\eta$ and the properties of the Kummer extensions in Proposition II.7. If $j\in\mathcal{S}^{\prime\prime}_{1}$, then any place lying over $P_{j}$ has degree two.
On the other hand, we have $N(E^{\prime}_{i})=2N^{\prime\prime}_{0}(\mathbf{s}_{i})-|\mathcal{T}(z_{i})|$ for $1\leq i\leq q^{d-1}-1$. From the Serre bound, we have the estimation for the number of rational places over the curve $E^{\prime}_{i}$ and it is given by
$\begin{split}|N(E^{\prime}_{i})-q-1|&\leq g(E^{\prime}_{i})\cdot\lfloor 2\sqrt{q}\rfloor\\ &\leq(d+1)\cdot\lfloor 2\sqrt{q}\rfloor.\end{split}$
Then we have the following estimation
$\begin{split}|N^{\prime\prime}_{1}(\mathbf{s}_{i})-N^{\prime\prime}_{0}(\mathbf{s}_{i})|&=\left|2N_{0}^{\prime\prime}-(q+1+t)\right|\\ &=|N(E_{i}^{\prime})+|\mathcal{T}(z_{i})|-(q+1+t)|\\ &\leq|N(E^{\prime}_{i})-q-1|+|t|+|\mathcal{T}(z_{i})|\\ &\leq(d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+d.\end{split}$
for any $1\leq i\leq q^{d-1}-1$. Therefore we have the desired results. ∎
Based on the results above, we have the following upper bound for the correlation.
Theorem III.6.
Let $q=p^{n}$ for odd prime $p$ and a positive integer $n$. Let $E/\mathbb{F}_{q}$ be a cyclic elliptic function field defined over $\mathbb{F}_{q}$ with $q+1+t$ rational places. Let $F$ be the fixed subfield of $E$ with respect to its translation group $T_{E}$ generated by $\sigma_{P}$. Let $Q$ be a place of $E$ with $\deg(Q)=d\geq 2$ such that $\gcd(d,q+1+t)=1$. Then the family of binary sequences $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ constructed above are with length $q+1+t$, size $q^{d-1}-1$ and correlation upper bounded by
$\begin{split}\mathrm{Cor}(\mathcal{S})&\leq(2d+1)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+2d.\end{split}$
Proof.
Let $\tau=\sigma_{P}$ be the generator of the translation group $T_{E}$. We begin by computing the autocorrelation of the family of the binary sequences $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ as in [9]. For each $1\leq i\leq q^{d-1}-1$, the autocorrelation of $\mathbf{s}_{i}$ at delay $u$ with $1\leq u\leq q+t$ is given by
[…]
The third equality follows from Lemma II.1. Since $z_{i}\in V\setminus\{0\}$ and $1\leq u\leq q+t$, we have $(z_{i}\tau^{-u}(z_{i}))_{\infty}=Q+\tau^{-u}(Q)$ by Proposition III.1. Consider the Kummer extension $E_{i}=E(y)$ with the defining equation
$E_{i}:y^{2}=z_{i}\cdot\tau^{-u}(z_{i}).$
Denote by $\mathcal{Z}(z_{i})$ the zeros of the rational function $z_{i}$ for $1\leq i\leq q^{d-1}-1$. By Proposition II.1 and Lemma 1, we know that places $Q$, $\tau^{-u}(Q)$, and zeros $\mathcal{Z}(z_{i})$ are all possible ramified places in $E_{i}/E$. From the Hurwitz genus formula, we have
$2g(E_{i})-2\leq 2(2g(E)-2)+\deg(Q)+\deg(\tau^{-u}(Q))+2\sum_{P\in\mathcal{Z}(z_{i})}\deg(P).$
Then the genus of $E_{i}$ satisfies
$g(E_{i})\leq 2d+1$
for each $1\leq i\leq q^{d-1}-1$.
Denote by $N_{0}$ the cardinality of the set $\mathcal{S}_{0}=\{0\leq j\leq q+t:\eta((z_{i}\tau^{-u}(z_{i})(P_{j}))=0\}$ and $N_{1}$ the cardinality of the set $\mathcal{S}_{1}=\{0\leq j\leq q+t:\eta((z_{i}\tau^{-u}(z_{i})(P_{j}))=1\}$. It is clear that
$N_{0}+N_{1}=q+1+t.$
If $j\in\mathcal{S}_{0}$, then $P_{j}$ splits into 2 rational places in $E_{i}$ or there is only one rational place lying over $P_{j}$ in $E_{i}$ by the definition of the quadratic residue map $\eta$ and the properties of the Kummer extensions in Proposition II.7. If $j\in\mathcal{S}_{1}$, then any place lying over $P_{j}$ has degree two.
Denote by $\mathcal{T}(z_{i})$ the set of zeros of $z_{i}$ which are rational, then the number of the rational points of $E_{i}$ can be given by $N(E_{i})=2N_{0}-|\mathcal{T}(z_{i}\tau^{-u}(z_{i}))|$ for each $1\leq i\leq q^{d-1}-1$.
From the Serre bound, we have
$\begin{split}|N(E_{i})-q-1|&\leq g(E_{i})\cdot\lfloor 2\sqrt{q}\rfloor\\ &\leq\left(2d+1\right)\cdot\lfloor 2\sqrt{q}\rfloor.\end{split}$
Therefore, we obtain the upper bound for the auto-correlation
[…]
The last inequality holds for the fact that $z_{i}\in\mathcal{L}(Q)$, then we have $|\mathcal{T}(z_{i})|\leq|\mathcal{Z}(z_{i})|\leq d$ for each $1\leq i\leq q^{d-1}-1$.
Now we consider the cross-correlation of the family of sequences. For the two distinct sequences $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ in $\mathcal{S}$ for $1\leq i\neq j\leq q^{d-1}-1$, the cross-correlation of $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ at delay $u$ with $0\leq u\leq q+t$ is given by
$\begin{split}\mathrm{C}_{u}(\mathbf{s}_{i},\mathbf{s}_{j})&=\sum^{q+t}_{k=0}(-1)^{\eta(z_{i}(P_{k}))+\eta(z_{j+u}(P_{k}))}\\ &=\sum^{q+t}_{k=0}(-1)^{\eta(z_{i}\tau^{-u}(z_{j})(P_{k}))}.\end{split}$
Consider the Kummer extension $E_{i,j}=E(y)$ with the defining equation
$E_{i,j}:y^{2}=z_{i}\cdot\tau^{-u}(z_{j}).$
If $1\leq u\leq q+t$, then we have $(z_{i}\tau^{-u}(z_{j}))_{\infty}=Q+\tau^{-u}(Q)$ by Proposition II.1. From the Hurwitz genus formula, we have
$\begin{split}2g(E_{i,j})-2&\leq 2(2g(E)-2)+\deg(Q)+\deg(\tau^{-u}(Q))+\sum_{P\in\mathcal{Z}(z_{i})}\deg(P)+\sum_{P\in\mathcal{Z}(z_{j})}\deg(P)\\ &\leq\deg(Q)+\deg(\tau^{-u}(Q))+2d=4d.\end{split}$
Hence, the genus of $E_{i,j}$ satisfies
$g(E_{i,j})\leq 2d+1$
for $1\leq i\neq j\leq q^{d-1}-1$. If $u=0$, then we have $(z_{i}\tau^{-u}(z_{j}))_{\infty}=(z_{i}z_{j})_{\infty}=2Q$ by Proposition III.1 and the genus of $E_{i,j}$ is also upper-bounded by $g(E_{i,j})\leq 2d+1$.
Let $N^{\prime}_{0}$ denote the cardinality of the set $\mathcal{S}_{0}^{\prime}=\{0\leq k\leq q+t:\eta(z_{i}\tau^{-u}(z_{j})(P_{k}))=0\}$ and $N^{\prime}_{1}$ denote the $\mathcal{S}^{\prime}_{1}=\{0\leq k\leq q+t:\eta(z_{i}\tau^{-u}(z_{j})(P_{k}))=1\}$. It is clear that
$N^{\prime}_{0}+N^{\prime}_{1}=q+1+t.$
If $j\in\mathcal{S}^{\prime}_{0}$, then $P_{j}$ splits into 2 rational places in $E_{i,j}$ or there is only one rational place lying over $P_{j}$ in $E_{i,j}$ by the definition of quadratic residue map $\eta$ and the properties of the Kummer extensions in Proposition II.7. If $j\in\mathcal{S}^{\prime}_{1}$, then any place lying over $P_{j}$ has degree two.
On the other hand, $N(E_{i,j})=2N^{\prime}_{0}-|\mathcal{T}(z_{i}\tau^{-u}(z_{j}))|$ for $1\leq u\leq q+t$. From the Serre bound, we have the estimation for the number of the rational places on the curve $E_{i,j}$
$\begin{split}|N(E_{i,j})-q-1|&\leq g(E_{i,j})\cdot\lfloor 2\sqrt{q}\rfloor\\ &\leq(2d+1)\cdot\lfloor 2\sqrt{q}\rfloor.\end{split}$
Therefore, we obtain the following upper bound for the cross-correlation
$\begin{split}|\mathrm{C}_{u}(\mathbf{s}_{i},\mathbf{s}_{j})|&=|N^{\prime}_{0}-N^{\prime}_{1}|=|2N^{\prime}_{0}-q-1-t|\\ &\leq|N(E_{i,j})-q-1|+|\mathcal{T}(z_{i}\tau^{-u}(z_{j}))|+|t|\\ &\leq\left(2d+1\right)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+|\mathcal{T}(z_{i}\tau^{-u}(z_{j}))|\\ &\leq\left(2d+1\right)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+|\mathcal{T}(z_{i})|+|\mathcal{T}(z_{j})|\\ &\leq\left(2d+1\right)\cdot\lfloor 2\sqrt{q}\rfloor+|t|+2d\end{split}$
for all $0\leq u\leq q+t$.
Combining the results of $|\mathrm{A}_{u}(\mathbf{s}_{i})|$ and $|\mathrm{C}_{u}(\mathbf{s}_{i},\mathbf{s}_{j})|$ for $1\leq i\neq j\leq q^{d-1}-1$, we have the desired results. ∎
Corollary III.7.
Let $n$ be an odd positive integer, $q=3^{n}$ and $t=\pm 3^{(n+1)/2}$. Then there exists a family of binary sequences with length $q+1+t$, size q1q-1 and the correlation upper-bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 3^{n/2}\rfloor+3^{(n+1)/2}+4$.
Proof.
For $t=\pm 3^{(n+1)/2}$, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ as $\gcd(2,3^{n}+1+t)=1$. Thus, the corollary follows from Theorem III.6. ∎
Corollary III.8.
Let $q=5^{n}$ for a positive integer $n$. Let $t=\pm 5^{n/2}$ and $n$ be even. Then there exists a family of binary sequences with length $q+1+t$, size q1q-1 and correlation upper bounded by $\mathrm{Cor}(\mathcal{S})\leq 5\cdot\lfloor 2\cdot 5^{n}\rfloor+|t|+4=11\cdot 5^{n/2}+4$.
Proof.
For $t=\pm 5^{n/2}$, there is a cyclic elliptic function field with $q+1+t$ rational places according to Lemma II.4. By Lemma II.5, there exists a place $Q$ in $E^{T_{E}}$ of degree $\deg(Q)=2$ that splits completely in $E$ for the fact that $\gcd(2,5^{n}+1+t)=1$. Thus, the corollary follows from Theorem III.6. ∎
III-C The Linear Complexity
In this subsection, we shall show that the family of binary sequences $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ constructed above also has a large linear complexity.
Theorem III.9.
The linear complexity of the family of binary sequences $\mathcal{S}=\{\mathbf{s}_{i}:1\leq i\leq q^{d-1}-1\}$ satisfies
$\mathrm{LC}(\mathcal{S})\geq\frac{q+1+2t-d-(d+1)\cdot\lfloor 2\sqrt{q}\rfloor}{d+d\cdot\lfloor 2\sqrt{q}\rfloor}.$
Proof.
We only need to prove that
$\ell(\mathbf{s}_{i})\geq\frac{q+1+2t-d-(d+1)\cdot\lfloor 2\sqrt{q}\rfloor}{d+d\cdot\lfloor 2\sqrt{q}\rfloor}$
for each sequence $\mathbf{s}_{i}\in\mathcal{S},1\leq i\leq q^{d-1}-1$.
Denote by $\ell=\ell(\mathbf{s}_{i})$. Then there exist $\ell+1$ elements such that $\lambda_{0}=\lambda_{\ell+1}=1$ and $\lambda_{2},\cdots,\lambda_{\ell}\in\mathbb{F}_{2}$ such that
$\sum^{\ell}_{j=0}\lambda_{j}\cdot\eta(z_{i}(P_{j+u}))=0$
for any integer $u\geq 1$. Let $\tau=\sigma_{P}$ be a generator of the translation group $T_{E}$. By the definition of linear complexity, we have
[…]
where $w_{i}=\prod^{\ell}_{j=0}(\tau^{-j}(z_{i}))^{\lambda_{j}}$. Hence, we obtain the vector
$(\eta(w_{i}(P_{1})),\eta(w_{i}(P_{2})),\cdots,\eta(w_{i}(P_{q+1+t})))=\mathbf{0}.$
Consider the Kummer extension $L_{i}/E$ with the defining equation
$L_{i}:y^{2}=\prod^{\ell}_{j=0}(\tau^{-j}(z_{i}))^{\lambda_{j}}=w_{i}.$
By Proposition III.1 and Lemma 1, we have $(w_{i})_{\infty}=\sum^{\ell}_{j=0}\lambda_{j}\tau^{-j}(Q)$ and the genus can also be calculated by
$\begin{split}2g(L_{i})-2&=2\cdot(2g(E)-2)+\deg(\mathrm{Diff}(L_{i}/E))\\ &\leq(\ell+1)\left(d+\sum_{P\in\mathcal{Z}(z_{i})}\deg(P)\right)\\ &\leq(\ell+1)\cdot 2d.\end{split}$
Then we have an upper bound for the genus $g(L_{i})$:
$g(L_{i})\leq(\ell+1)d+1.$
It follows that
$2(q+1+t)-|\mathcal{T}(w_{i})|\leq N(L_{i})\leq q+1+\left[(\ell+1)d+1\right]\cdot\lfloor 2\sqrt{q}\rfloor.$
where $\mathcal{T}(w_{i})$ denotes the set of zeros that are rational of $w_{i}$ for $1\leq i\leq q^{d-1}-1$. Since $|\mathcal{T}(w_{i})|\leq|\mathcal{Z}(w_{i})|$, we have
$2(q+1+t)-|\mathcal{Z}(w_{i})|\leq N(L_{i})\leq q+1+\left[(\ell+1)d+1\right]\cdot\lfloor 2\sqrt{q}\rfloor.$
Furthermore, we have the following inequalities
$\begin{split}2(q+1+t)&\leq|\mathcal{Z}(w_{i})|+q+1+\left[(\ell+1)d+1\right]\cdot\lfloor 2\sqrt{q}\rfloor\\ &\leq(\ell+1)\cdot\left(d+d\cdot\lfloor 2\sqrt{q}\rfloor\right)+q+1+\lfloor 2\sqrt{q}\rfloor.\end{split}$
Then we have
$\ell(\mathbf{s}_{i})\geq\frac{q+1+2t-d-(d+1)\cdot\lfloor 2\sqrt{q}\rfloor}{d+d\cdot\lfloor 2\sqrt{q}\rfloor}\\$
for all $1\leq i\leq q^{d-1}-1$ and we have the desired result. ∎
III-D Numerical Results
The previous subsections provide a theoretical framework for constructing a new family of binary sequences with low correlation and high linear complexity by utilizing cyclic elliptic curves. In this subsection, we provide some numerical results of our construction of binary sequences. It turns out that the actual correlation of our family of binary sequences is significantly lower than the theoretical upper bound provided in Theorem III.6.
In the following, we also consider the cases for the place $\deg(Q)=d=2$ and the cyclic elliptic function fields over the finite fields with characteristic $p=3$.
First, consider the balance. Let $q=3^{n}$. It is also obvious that $\gcd(2,3^{n})=1$. By Corollary II.4, there exists a cyclic elliptic function field $E/\mathbb{F}_{q}$ with $q$ rational places. By Theorem III.5, we can construct a family of binary sequences with size $3^{n}-1$, length $q$ and absolute difference $\Delta(\mathcal{S})$ upper bounded by $3\cdot\lfloor 2\sqrt{q}\rfloor+2$. We list some numerical results in table II and it turns out that the actual value $|N^{\prime\prime}_{1}(\mathbf{s})-N^{\prime\prime}_{0}(\mathbf{s})|$ for the sequence $\mathbf{s}\in\mathcal{S}$ is much smaller than the theoretical bound given in (2).
[…]
Field Size Family Size $N^{\prime\prime}_{0}(\mathbf{s})$ $N^{\prime\prime}_{1}(\mathbf{s})$ $|N_{1}^{\prime\prime}(\mathbf{s})-N_{0}^{\prime\prime}(\mathbf{s})|$
81 80 4063 4037 26
243 242 5991 6159 168
729 728 18302 18148 154
Now we consider the actual correlation. Also let the field size be $q=3^{n}$. By the results in Theorem I, we can construct a family of binary sequences with size $3^{n}-1$, length $q$ and correlation upper bounded by $5\cdot\lfloor 2\sqrt{q}\rfloor+4$. We list some results calculated by SageMath [2] for this case in Table III.
[…]
Field Size Sequence Length Family Size Theoretical Bound of Correlation Actual Correlation
81 81 80 94 43
243 243 242 159 73
729 729 728 274 131
2187 2187 2186 469 274
Now consider another construction of the binary sequences from the cyclic elliptic function fields over $\mathbb{F}_{q}$. Let $t=\sqrt{q}$ if $n$ is even, or $\sqrt{3q}$ if $n$ is odd. By Corollary II.4, we can construct a family of binary sequences with family size $3^{n}-1$, length $q+1+t$ and the correlation upper bounded by $5\cdot\lfloor 2\sqrt{q}\rfloor+t+4$. We list some numerical results in Table IV.
[…]
Field Size Sequence Length Family Size Theoretical Bound of Correlation Actual Correlation
81 91 80 103 46
243 271 242 186 78
729 757 728 251 148
For the detailed programming code, we list it on the webpage https://github.com/lxfhah-byte/binary-sequences-odd.
IV Conclusions and Future Work
In this paper, we extend the construction of binary sequences given by Jin et al. in [9] via the cyclic elliptic function fields with characteristic $2$ to general odd characteristics by applying the quadratic residue map $\eta$ and the theory of Kummer extensions of function fields. Since our construction is over elliptic function fields with odd characteristic, our constructions have flexible parameters. From the numerical experiments, it is noticed that the real correlation is much better than the theoretical upper bound given in Theorem III.6. In particular, for applications in cryptography, we also investigate the property of balance for this family of binary sequences. It turns out that the actual bound of $\Delta(\mathcal{S})$ is also much smaller than the theoretical bound. For the future work, it is interesting to 1) give better bounds for correlations and linear complexities of the sequences constructed in [9] and this paper by using other deep mathematical methods and 2) generalize the constructions of binary sequences by utilizing general algebraic function fields.